Mr. Ayala borrows 100,000 pesos at 10% effective annual interest. He must pay back the loan over 30 years with uniform monthly payments due on the first day of each month. What does Mr. Ayala pay each month?
Answer #1

Calculation of interest rate per month

i+1 = (1 + I/12 )12

1.10 = (1+ I /12 )12

I /12 = 0.007974

Total number of payments = 12 months *30 years =360 months

Borrowed money = A + P

100,000 = A + A(1+0.007974)359 - 1 /(1+0.007974)359 0.007974

A= 839.19 each month.